# Beyond the Use-and-then-Forget (UatF) Bound: Fixed Point Algorithms for Statistical Max-Min Power Control

### v2

@@ IV Simulations
We consider a cell-less network with $L=16$ access points, each equipped with a four-element uniform linear array with half-wavelength spacing ($\lambda/2$), where $\lambda=c/f_{\mathrm{c}}$ is the wavelength, $f_{\mathrm{c}}$ is the carrier frequency, and $c$ is the propagation speed. The access points are placed uniformly over a $1000\times 1000~\mathrm{m}^{2}$ area. We distribute $N=25$ single-antenna users uniformly at random and apply a conventional wrap-around technique to emulate an unbounded service area. The user-access point antenna height difference is $\Delta h=11$ m. The large-scale fading (path loss) $\beta_{u,a}$ between user $u\in\mathcal{U}$ and access point $a\in\mathcal{A}$ follows the COST-231 Walfish–Ikegami UMi model [23, Sect. 5.2]:
@@
where $d_{u,a}$ is the 3D distance and $F_{u,a}\sim\mathcal{N}(0,\sigma_{\mathrm{sf}}^{2})$ models shadow fading (in dB). Small-scale fading follows a spatially correlated Rician model. All beamformers operate on channel estimates – rather than the true channels – obtained using the estimation procedure described in [7]. For pilot assignment and formation of user-centric cooperation clusters in the cell-less network, we use the algorithm described in [2, Algorithm 4.1]. The main simulation parameters are listed in Table I, and we refer the readers to [7] for further details on the simulation.
@@
The performance evaluation considers uniform weights in (12) (i.e., $(\forall u\in\mathcal{U})~\alpha_{u}=1$) and centralized cell-less MMSE beamforming design with fixed (unitary) power weighting coefficients in the channel inversion stage [9, Sect. 3-D]. We estimate the ergodic rates by averaging across 500 Monte Carlo trials, each drawing an independent channel realization for a single, fixed network layout. In the proposed fixed point iterations, shown in Corollary 1, we use (5) as the utility function, which is called “optimistic ergodic rate” (OER) in the following (see Remark 3). As a baseline, we compute the OER bound with the solution to the standard max-min problem under the UatF bound, which admits a pseudo-closed-form expression [9, Proposition 4]. This approach mirrors established practice in the literature: if simple fixed point algorithms capable of solving (12) were unknown (as was the case prior to this study, to the best of our knowledge), we would consider the surrogate max-min power-control problem that uses the UatF bound instead of the target bound in (5), with the expectation that the surrogate solution would approximate the solution to (12).

### v3

@@ IV Simulations
We consider a cell-less network with $L=16$ access points, each equipped with a four-element uniform linear array with half-wavelength spacing ($\lambda/2$), where $\lambda=c/f_{\mathrm{c}}$ is the wavelength, $f_{\mathrm{c}}$ is the carrier frequency, and $c$ is the propagation speed. The access points are placed uniformly over a $1000\times 1000~\mathrm{m}^{2}$ area. We distribute $N=25$ single-antenna users uniformly at random and apply a conventional wrap-around technique to emulate an unbounded service area. The user-access point antenna height difference is $\Delta h=11$ m. The large-scale fading (path loss) $\beta_{u,a}$ between user $u\in\mathcal{U}$ and access point $a\in\mathcal{A}$ follows the COST-231 Walfish–Ikegami UMi model [1, Sect. 5.2]:
@@
where $d_{u,a}$ is the 3D distance and $F_{u,a}\sim\mathcal{N}(0,\sigma_{\mathrm{sf}}^{2})$ models shadow fading (in dB). Small-scale fading follows a spatially correlated Rician model. All beamformers operate on channel estimates – rather than the true channels – obtained using the estimation procedure described in [2]. For pilot assignment and formation of user-centric cooperation clusters in the cell-less network, we use the algorithm described in [11, Algorithm 4.1]. The main simulation parameters are listed in Table I, and we refer the readers to [2] for further details on the simulation.
@@
The performance evaluation considers uniform weights in (12) (i.e., $(\forall u\in\mathcal{U})~\alpha_{u}=1$) and centralized cell-less MMSE beamforming design with fixed (unitary) power weighting coefficients in the channel inversion stage [20, Sect. 3-D]. We estimate the ergodic rates by averaging across 500 Monte Carlo trials, each drawing an independent channel realization for a single, fixed network layout. In the proposed fixed point iterations, shown in Corollary 1, we use (5) as the utility function, which is called “optimistic ergodic rate” (OER) in the following (see Remark 3). As a baseline, we compute the OER bound with the solution to the standard max-min problem under the UatF bound, which admits a pseudo-closed-form expression [20, Proposition 4]. This baseline mirrors established practice in the literature: if simple fixed point algorithms capable of solving (12) were unknown (as was the case prior to this study, to the best of our knowledge), we would consider the surrogate max-min power-control problem that uses the UatF bound instead of the target bound in (5), with the expectation that the surrogate solution would approximate the solution to (12).